OpenStudy (anonymous):

Given the system of equations, what is the solution?2x + y = -1x - y = -5

OpenStudy (anonymous):

x=-2 y=3

OpenStudy (anonymous):

2x+y = -1 x-y = -5 --------- 3x=-6 x=-2 x-y=-5 -2-y=-5 -y=-3 y=3

OpenStudy (anonymous):

thanks yoloswagoo!

OpenStudy (anonymous):

anytime! buddeh! :3